OpenStudy (anonymous):

In the game of blackjack, a 2-card hand consisting of an ace and either a face card or a 10 is called a "blackjack." If a standard 52-card deck is used, determine how many blackjack hands can be dealt. (A "face card" is a jack, queen, or king.)

OpenStudy (kinggeorge):

Well, first off, how many ace cards are there, and how many face cards are there?

OpenStudy (kinggeorge):

Once you find how many ace cards and how many face cards there are, just multiply them together.

OpenStudy (anonymous):

there are 4 ace cards and 4 of each face card and 4 10s so do i mulitiple \[4\times4^{4}\times4\]

OpenStudy (kinggeorge):

There are definitely 4 aces, 4 of each face card, and 4 tens. If you combine the face cards with the tens, you have 4 face cards or tens of each suit, so there are \(4*4=16\) total face cards or tens. So there should be \(4*16 = 4^3=64\) blackjack hands.

OpenStudy (anonymous):

kk thanks

OpenStudy (kinggeorge):

You're welcome.